What is likely to be the principal quantum number for a circular orbit of diameter 20nm of the H atom if we assume Bhor orbit to be the same as that represented by the principal quantam number?

Grade:11

3 Answers

mohit sengar
13 years ago

radius of bohr's orbit is  10 nm 

also, radius of bohr's orbit is given by  0.0529 ×n²/z  nm  as z=1  , therefore  r= 0.0529n²  nm 

equating the above two ,we have  

           10 = 0.0529 n²

on solving this u get principal quantumm no = 13 

Sneha Das
5 years ago
Given: diameter = 20nm
thus, radius=10nm
r of n th shell= 0.539 X n X n
1nm=0.1A    thus, 0.539A=0.0539nm
thus, 10=0.0539 x n x n
n x n =185.528
n=13.62
orbit can’​t be 13.62 thus it is 14
Anand singh
5 years ago
10Å=1nm
Given diameter is 20nm
radius = 10nm=100Å
z=1
Radius =0.529×n²/z
100/0.529=n²
n²≈100
n=10
So orbit number is 10